# A symmetry-based approach to species-rich ecological communities

### v1

@@ I Introduction
Most theoretical work so far has revolved around the so-called random generalized Lotka-Volterra model (gLV), where interactions are sampled independently from each other [3]. Despite recent qualitative agreement with some of the predictions of these models [7], many features of ecological communities remain elusive to ‘disordered ecology’[8]. Whether these mismatches are the signature of non-random structure in the interactions, and what elements should be added to the models to make them more realistic, are important research questions that have fostered developments in several directions [9, 10, 11, 12].
@@ II.1 The Lotka-Volterra equations
We consider a set of $S$ interacting species, labeled $i=1,\dots,S$ and model the temporal dynamics of their respective abundances, which we denote $x_{i}(t)$, according to the generalized Lotka-Volterra equations (gLV),
@@
Apart from its interaction with the environment, the net growth rate of the species depends on its interaction with the rest of the community, as represented by the matrix of interaction coefficients $A_{ij}$. This matrix encodes, in an abstract form, the information about all ecological processes that are relevant to understand the dynamics. As such, specifying the special characteristics of a community is tantamount to defining an interaction matrix. Tthe interaction matrix alone accounts for $S^{2}$ parameters. For $S\gg 1$, this is a prohibitively large number. Reducing the number of parameters is critical to recover a useful and tractable model.
@@ II.2 Our definition of structure
Once randomization is done, the only remaining information about the original interaction matrix are its eigenvalues, which are common to the matrices of the ensemble thus defined. In particular, the spectrum encodes how the interactions were originally constructed.
In the simplest case, the spectrum is semi-circular, and the ensemble considered is the Gaussian Orthogonal Ensemble (GOE), where the $A_{ij}$ are (up to the symmetry constraint) independent and identically distributed Gaussian variables. For other spectral distributions, they display a diffuse correlation pattern, and are not Gaussian when considered as a whole. In the following section, we show how these diffuse correlation patterns have an important impact in the dynamics.
@@ III.1 General setting
where $\mu$ stands for the Haar measure over the orthogonal group, $p$ is the distribution of the initial conditions (which might be deterministic or random) and $\delta$ is a functional Dirac delta that enforces the trajectory to undergo the prescribed dynamics. The goal is then to simplify this expression up to a point where the probability distribution $\mathrm{d}\mathbb{P}\left[x_{i}(t)\right]$ can be recognized explicitly in a simple form. The strategy to do so is standard in disordered systems[23, 4], hence we mainly focus on the results, and relegate details to SI (see SI Sec. LABEL:SI-sec:dmft). The key step, performing the integral over $U$, relies on the following large deviation principle for the so-called HCIZ integral [24, 25],
@@
$\frac{\mathrm{d}x_{i}}{\mathrm{d}t}=x_{i}\left[1-ux_{i}+\int_{0}^{t}H(t,s)x_{i}(s)+\xi_{i}(t)\right],$ (8)
@@ III.2 Stationary regimes
$\begin{split}\hat{H}(\omega)&=G_{\rho}^{\prime}\left(\hat{K}(\omega)\right)\\ \hat{C_{\xi}}(\omega)&=\hat{C}(\omega)\frac{\mathrm{Im}\left[G_{\rho}^{\prime}\left(\hat{K}(\omega)\right)\right]}{\mathrm{Im}\left[\hat{K}(\omega)\right]}\end{split}$ (9)
@@ III.3 Fixed points
$\begin{split}x(t)&\rightarrow x\\ \xi(t)&\rightarrow\xi\\ C(t,t)&\rightarrow q\\ C_{\xi}(t,t)&\rightarrow z\\ \int_{0}^{t}\mathrm{d}sK(t,s)&\rightarrow\chi\\ \int_{0}^{t}\mathrm{d}sH(t,s)&\rightarrow\eta\end{split}$ (10)
@@
$\begin{split}\eta&=G^{\prime}_{\rho}\left(\chi\right)\\ z&=qG^{\prime\prime}_{\rho}\left(\chi\right)\\ q&=\langle x^{2}\rangle=\frac{z}{(u-\eta)^{2}}\omega_{2}\\ \chi&=\frac{\omega_{0}}{u-\eta}\end{split}$ (12)
@@
Known results can be obtained by specializing $\rho$ to previously studied distributions. For instance, if $\rho$ is the semi-circle distribution with variance $\sigma^{2}$ then $G_{\rho}(x)=\sigma^{2}x^{2}/2$ and one obtains the well-known equations for the GOE case [5, 3, 4]. For other spectral distributions, $G_{\rho}$ contains higher-order cumulants that influence equilibrium properties. These are the signature of correlations between interactions at the level of the whole matrix, the GOE case being the very special case where those correlations are identically zero.
@@
Fig. 1 shows the Species Abundance Distribution (SAD) for these three classes of spectra. Importantly, the spectra were tailored so as to all have the same variance. The differences in the results are therefore due to the higher-order cumulants encoded in $G_{\rho}$ or, in other terms, to the different shapes of each spectrum. By considering only the variance of the interactions and neglecting the shape of the spectra, a random model based on the GOE ensemble would miss these differences.
@@
Figure 1: Influence of the spectral distribution on equilibrium abundances: The Species Abundance Distribution (SAD) is plotted for three different spectra. Blue: A shifted and scaled Marcenko-Pastur distribution, corresponding to Hebbian interactions. Orange: A uniform distribution. Green: A binary distribution, with only two unique values. The spectral distributions all have a variance of $0.4$ and the self-regulation is set to $u=1$, but the ensuing SADs are notably different.
@@ IV.2 Linear instability
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\delta x_{i}&=x_{i}^{\infty}\Big{[}-(u-\eta)\delta x_{i}(t)+\delta\xi_{i}(t)\Big{]}\\ &+\delta x_{i}(t)\Big{[}1-(u-\eta)x_{i}^{\infty}+\xi_{i}^{\infty}\Big{]},\end{split}$ (14)
@@
$\left(u-\eta\right)^{2}-\omega_{0}G_{\rho}^{\prime\prime}\left(\frac{\omega_{0}}{u-\eta}\right)=0$ (16)
@@
$\lim_{x\rightarrow\infty}x^{2}G_{\rho}^{\prime\prime}(x)>\frac{1}{2}$ (18)
@@ Appendix B The function $G_{\rho}$
where $\kappa_{n}$ are the so-called free cumulants of $\rho$. For our needs the existence of the cumulants is guaranteed by the fact that we’ll only consider cases where $\rho$ is bounded. In particular, one may take the previous series expansion and use it to extend the definition or $R$ and $G$ to the complex plane. For the examples of the main text this gives
@@
Uniform $\frac{\sqrt{\epsilon}}{\tanh(\sqrt{\epsilon}x)}-\frac{1}{x}$ $\frac{1}{x^{2}}-\frac{\epsilon}{\sinh^{2}(\sqrt{\epsilon}x)}$
Binary $\frac{\sqrt{1+4\lambda x^{2}}-1}{2x}$ $\frac{\sqrt{1+4\lambda x^{2}}-1}{2x^{2}\sqrt{1+4\lambda x^{2}}}$

### v2

@@ I Introduction
Most theoretical work so far has revolved around the so-called random generalized Lotka-Volterra model (gLV), where interactions are sampled independently from each other [3]. Despite recent qualitative agreement with some of the predictions of these models [7], many features of ecological communities remain elusive to ‘disordered ecology’[8]. Whether these mismatches are the signature of non-random structure in the interactions, and what elements should be added to the models to make them more realistic, are important research questions that have fostered developments in several directions [9, 10, 11, 12, 13].
@@ II.1 The Lotka-Volterra equations
We consider a set of $S$ interacting species, labeled by $i=1,\dots,S$ and model the temporal dynamics of their respective abundances, which we denote $x_{i}(t)$, according to the generalized Lotka-Volterra equations (gLV),
@@
Apart from its interaction with the environment, the net growth rate of the species depends on its interaction with the rest of the community, as represented by the matrix of interaction coefficients $A_{ij}$. This matrix encodes, in an abstract form, the information about all ecological processes that are relevant to understand the dynamics. As such, specifying the special characteristics of a community is tantamount to defining an interaction matrix. The interaction matrix alone accounts for $S^{2}$ parameters. For $S\gg 1$, this is a prohibitively large number. Reducing the number of parameters is critical to recover a useful and tractable model.
@@ II.2 Our definition of structure
Once randomization is done, the only remaining information about the original interaction matrix are its eigenvalues, which are common to the matrices of the ensemble thus defined. In particular, the spectrum encodes how the interactions were originally constructed. Importantly, the second moment of the spectral distribution determines the variance of the interaction coefficients, which equals $\langle\lambda^{2}\rangle/S$.
In the simplest case, the spectrum is semi-circular, and the ensemble considered is the Gaussian Orthogonal Ensemble (GOE), where the $\left(A_{ij}\right)_{i<j}$ are jointly distributed as independent Gaussians, and $k$-point statistics of the form $\langle A_{i_{1}j_{1}}\dots A_{i_{k}j_{k}}\rangle$ vanish at any order $k$. For other spectral distributions, we show in Appendix C that such statistics still vanish asymptotically for $k\ll S$, but don’t when the number of terms is comparable to the size of the system. Thus coefficients appear to be independently Gaussian when taken in small sets but not when considering the whole matrix. This diffuse pattern of high-order inter-dependencies thus has an impact when considering quantities that depend on an extensive number of interaction coefficients, such as the terms $\sum_{j}A_{ij}x_{j}$ in (1). Thus they have an impact on the dynamics, as we show next.
@@ III.1 General setting
where $\mu$ stands for the Haar measure over the orthogonal group, $p$ is the distribution of the initial conditions (which might be deterministic or random) and $\delta$ is a functional Dirac delta that enforces the trajectory to undergo the prescribed dynamics. The goal is then to simplify this expression up to a point where the probability distribution $\mathrm{d}\mathbb{P}\left[x_{i}(t)\right]$ can be recognized explicitly in a simple form. The strategy to do so is standard in disordered systems[24, 4], hence we mainly focus on the results, and relegate details to SI (see SI Sec. LABEL:SI-sec:dmft). The key step, performing the integral over $U$, relies on the following large deviation principle for the so-called HCIZ integral (see [25, 26, 27, 28] as well as SI Sec. LABEL:SI-sec:free),
@@
$\frac{\mathrm{d}x_{i}}{\mathrm{d}t}=x_{i}\left[1-ux_{i}+\int_{0}^{t}H(t,s)x_{i}(s)+\xi_{i}(t)\right].$ (8)
@@ III.2 Stationary regimes
$\begin{split}\hat{H}(\omega)&=G_{\rho}^{\prime}\left(\hat{K}(\omega)\right)\\ \hat{C_{\xi}}(\omega)&=\hat{C}(\omega)\frac{\mathrm{Im}\left[G_{\rho}^{\prime}\left(\hat{K}(\omega)\right)\right]}{\mathrm{Im}\left[\hat{K}(\omega)\right]}.\end{split}$ (9)
@@ III.3 Fixed points
$\begin{split}x(t)&\rightarrow x\\ \xi(t)&\rightarrow\xi\\ C(t,t)&\rightarrow q\\ C_{\xi}(t,t)&\rightarrow z\\ \int_{0}^{t}\mathrm{d}sK(t,s)&\rightarrow\chi\\ \int_{0}^{t}\mathrm{d}sH(t,s)&\rightarrow\eta.\end{split}$ (10)
@@
$\begin{split}\eta&=G^{\prime}_{\rho}\left(\chi\right)\\ z&=qG^{\prime\prime}_{\rho}\left(\chi\right)\\ q&=\langle x^{2}\rangle=\frac{z}{(u-\eta)^{2}}\omega_{2}\\ \chi&=\frac{\omega_{0}}{u-\eta}.\end{split}$ (12)
@@
Known results can be obtained by specializing $\rho$ to previously studied distributions. For instance, if $\rho$ is the semi-circle distribution with variance $\sigma^{2}$ then $G_{\rho}(x)=\sigma^{2}x^{2}/2$ and one obtains the well-known equations for the GOE case [5, 3, 4]. For other spectral distributions, $G_{\rho}$ contains higher-order cumulants that influence equilibrium properties. In Sec. II.2 we identified such higher-order terms as a signature of inter-dependencies between interactions at the level of the whole matrix. Here, we see how they impact the dynamics at the collective level, through their influence on aggregated quantities such as $z$ and $\eta$.
@@
Fig. 1 shows the Species Abundance Distribution (SAD) for these three classes of spectra. Thus . Importantly, the spectra were tailored so as to all have the same variance. Therefore the variance in the interaction coefficients is the same for all three classes. The differences in the results are therefore due to the higher-order cumulants encoded in $G_{\rho}$ or, in other terms, to the different shapes of each spectrum. By considering only the variance of the interactions and neglecting the shape of the spectra, a random model based on the GOE ensemble would miss these differences.
@@
Figure 1: Influence of the spectral distribution on equilibrium abundances: The Species Abundance Distribution (SAD) is plotted for three different spectra (bars are simulations, lines are predictions). Blue: A shifted and scaled Marcenko-Pastur distribution, corresponding to Hebbian interactions. Orange: A uniform distribution. Green: A binary distribution, with only two unique values. The spectral distributions all have a variance of $0.4$ and the self-regulation is set to $u=1$, but the ensuing SADs are notably different.
@@ IV.2 Linear instability
$\begin{split}\frac{\mathrm{d}}{\mathrm{d}t}\delta x_{i}&=x_{i}^{\infty}\Big{[}-(u-\eta)\delta x_{i}(t)+\delta\xi_{i}(t)\Big{]}\\ &+\delta x_{i}(t)\Big{[}1-(u-\eta)x_{i}^{\infty}+\xi_{i}^{\infty}\Big{]}.\end{split}$ (14)
@@
$\left(u-\eta\right)^{2}-\omega_{0}G_{\rho}^{\prime\prime}\left(\frac{\omega_{0}}{u-\eta}\right)=0.$ (16)
@@
$\lim_{x\rightarrow\infty}x^{2}G_{\rho}^{\prime\prime}(x)>\frac{1}{2}.$ (18)
@@ Appendix B The function $G_{\rho}$
where $\kappa_{n}$ are the so-called free cumulants of $\rho$. For our needs the existence of the cumulants is guaranteed by the fact that we’ll only consider cases where $\rho$ is bounded. In particular, one may take the previous series expansion and use it to extend the definition or $R$ and $G$ to the complex plane.
The examples used in the main text are:
  1. 1.
    A Wigner distribution with variance $\sigma^{2}$.
  2. 2.
    A Marcenko-Pastur distribution with ratio $\alpha$.
  3. 3.
    A uniform distribution with edges ±$\pm\epsilon$.
  4. 4.
    A binary distribution with half the eigenvalues at $-\lambda$ and the other half at $\lambda$.
For these we have:
@@
Uniform $\frac{\epsilon}{\tanh(\epsilon x)}-\frac{1}{x}$ $\frac{1}{x^{2}}-\frac{\epsilon^{2}}{\sinh^{2}(\epsilon x)}$
Binary $\frac{\sqrt{1+4\lambda^{2}x^{2}}-1}{2x}$ $\frac{\sqrt{1+4\lambda^{2}x^{2}}-1}{2x^{2}\sqrt{1+4\lambda^{2}x^{2}}}$
Appendix C Relationship between $G_{\rho}$ and the interaction coefficients
The joint distribution of a set of interaction coefficients $A_{i_{1}j_{1}},\dots,A_{i_{k}j_{k}}$ can be characterized through their generating function
$f(\lambda_{1},\dots,\lambda_{k})=\left\langle\exp\left(\lambda_{1}A_{i_{1}j_{1}}+\dots+\lambda_{k}A_{i_{k}j_{k}}\right)\right\rangle,$
[…]
$f(\lambda_{1},\dots,\lambda_{k})\sim\exp\left(\frac{S}{2}\mathrm{Tr}G_{\rho}\left(\frac{B}{S}\right)\right).$
Assume for simplicity that the indices $i_{p},j_{p}$ are disjoint between pairs with different $p$. Then the eigenvalues of $B$ are the ±$\pm\lambda_{p}$, and thus
$\begin{split}f(\lambda_{1},\dots,\lambda_{k})&\sim\exp\left(\frac{S}{2}\sum_{p}G_{\rho}\left(\frac{\lambda_{p}}{S}\right)\right)\\ &\sim\exp\left(\frac{\lambda_{p}^{2}\sigma^{2}}{2S}+\mathcal{O}(S^{-3})\right),\end{split}$
where $\sigma^{2}=G_{\rho}^{\prime\prime}(0)=\langle\lambda^{2}\rangle$ is the variance of the spectrum. We recognize the generating function of a set of centered Gaussian variables with common variance $\sigma^{2}/S$ and two-point correlations of order at most $S^{-2}$. Thus, correlations between small sets of coefficients vanish.
Inter-dependencies between coefficients appear when an extensive number of them is considered at once. For example, let us study the sum of all interaction coefficients $\bra{1}A\ket{1}$. Its generating function is
$\begin{split}f(\lambda)=\left\langle\exp\Big{(}\lambda\bra{1}A\ket{1}\Big{)}\right\rangle&\sim\exp\left(\frac{S}{2}\mathrm{Tr}G_{\rho}\left(\frac{2\lambda\ket{1}\bra{1}}{S}\right)\right)\\ &\sim\exp\left(\frac{S}{2}G_{\rho}(2\lambda)\right),\end{split}$
and therefore depends on the shape of $G_{\rho}$. In particular, the sum of the coefficients is not Gaussian, showing that the coefficients themselves are not jointly Gaussian. A Gaussian generating function is only recovered when the spectrum is semi-circular and $G_{\rho}(x)=\sigma^{2}x^{2}/2$.